# Concept Corrector: Erase concepts on the fly for text-to-image diffusion models

### v1

@@ Abstract
Text-to-image diffusion models have demonstrated the underlying risk of generating various unwanted content, such as sexual elements. To address this issue, the task of concept erasure has been introduced, aiming to erase any undesired concepts that the models can generate. Previous methods, whether training-based or training-free, have primarily focused on the input side, i.e. texts. However, they often suffer from incomplete erasure due to limitations in the generalization from limited prompts to diverse image content. In this paper, motivated by the notion that concept erasure on the output side, i.e. generated images, may be more direct and effective, we propose to check concepts based on intermediate-generated images and correct them in the remainder of the generation process. Two key challenges are identified, i.e. determining the presence of target concepts during generation and replacing them on the fly. Leveraging the generation mechanism of diffusion models, we present the Concept Corrector, which incorporates the Generation Check Mechanism and the Concept Removal Attention. This method can identify the generated features associated with target concepts and replace them using pre-defined negative prompts, thereby achieving concept erasure. It requires no changes to model parameters and only relies on a given concept name and its replacement content. To the best of our knowledge, this is the first erasure method based on intermediate-generated images. The experiments on various concepts demonstrate its impressive erasure performance. Code: https://github.com/RichardSunnyMeng/ConceptCorrector.
Figure 1: In this paper, we present Concept Corrector, a training-free method to erase concepts based on intermediate-generated images during the text-to-image diffusion process. Concept Corrector integrates the Generation Check Mechanism alongside Concept Removal Attention, to ascertain the presence of concepts during the generation process and erase them in the subsequent generations.  *   indicates the masked sensitive regions for display purposes. SD: Stable Diffusion. Please refer to Appendix J for more visualizations.
@@ 1 Introduction
The task of concept erasure has been proposed to address this issue [42, 22, 12]. Its goal is to prevent models from generating content related to certain concepts. Intuitively speaking, if the generation path from the textual inputs to the target visual concepts can be cut off, the resultant images, being guided by those texts, would naturally exclude the presence of these concepts. Researchers have explored various methods based on this idea, which can be categorized into training-based and training-free methods. Within the area of training-based methods, CA [22] and ESD [12] fine-tune the generation distributions conditioned on the texts containing target concepts. Instead of gathering texts, Bui et al. [2] and Meng et al. [30] incorporate a learnable prompt in the training and other works [34, 21, 55, 18] introduce adversarial training to improve the robustness of concept erasure. UCE [13], MACE [29], and RECE [14] edit cross-attention weights by aligning the keys and values of target concepts to others. There are also some methods [10, 49, 4] ablating model parameters based on their sensitivity to related prompts. While these methods fine-tune diffusion models, Latent Guard [26] and GuardT2I [51] propose to fine-tune the text encoders. Within the area of training-free methods, SLD [42] uses the texts of concepts to steer generation in the opposite direction. SAFREE [52] maps the prompt embeddings to those of target concepts and removes the corresponding components.
@@
While most studies explore text-based concept erasure methods, little attention has been paid to accomplishing this task based on the output of generative models, i.e. generated images. Text-based erasure methods often face the challenge of prompt generalization [46, 56, 50, 6]. It arises due to the difficulty in using prompts, which are collected or learned by these methods, to comprehensively cover the diverse image content associated with target concepts. Consequently, the effectiveness of the erasure is limited. If the focus is shifted to images, the target of erasure becomes more directly addressed, potentially leading to a more effective erasure performance. A common practice is to incorporate a safety checker following the generation process to filter out unwanted content like in Stable Diffusion [37], which works well but cannot correct generated content to obtain images like text-based erasure methods.
In this paper, our objective is to actively intervene in the generation process. Through checking and correction of intermediate-generated images, we seek to erase any target concepts in the final images. To complete this task, we need to answer two questions. First, how to check the content of images that are still in the generation process? Second, how to locate and replace concept-related content in the subsequent generation process? For the first question, we exploit the generation mechanism of diffusion models and feed predicted outputs into a binary concept classifier to determine the presence of concepts. For the second question, we propose Concept Removal Attention, a variation of the existing cross attention mechanism, to correct concept-related features, thereby removing target concepts in the final generations. The above explorations form the method we present, namely Concept Corrector.
@@
Compared with previous methods, it has significant advantages in the erasure performance of various concepts. Firstly, this method checks the presence of concepts in images during generation, thereby providing a more direct assessment of concepts. Secondly, it corrects concept-related content in the subsequent generation process. As a result, it directly outputs images in which target concepts are erased. Moreover, similar to other training-free methods, all parameters remain unchanged, which saves the time of model training, protects model knowledge, and facilitates flexible adjustment of target concepts. In the experiments, we evaluate the erasure performance using user prompts and adversarial prompts. The evaluated concepts include Not-Safe-For-Work, objects, and painting styles. As shown in Fig.1, Tab.1 and Tab.2, our method achieves impressive erasure performance, significantly reducing the generation of most concepts to within 5% while other methods are still far from this level of performance. In summary, the contributions of this paper can be outlined as follows.
@@
  • We carefully analyze the feasibility of using intermediate images for checking concepts, and discuss the effect of localizing and correcting concepts using model parameters without relying on input prompts.
@@
  • The above explorations forms Concept Corrector, a novel method to erase concepts during the generation process. To our knowledge, this is the first work to achieve erasure based on intermediate-generated images.
@@
  • The experiments and visualizations demonstrate the impressive effectiveness of our method. A series of ablation experiments are conducted to discuss the conponents of the proposed method.
@@ 4.2 Generation Check Mechanism
As mentioned in Sec.3.1, the final generation results can be predicted at intermediate time steps. This is the foundation for our method to check concepts during the diffusion generation. Some previous studies [7, 19] further explore the generative traits of different diffusion stages. A common observation is that diffusion models initially generate global structures and then refine these with local details as the diffusion progresses. Building upon this insight, we set two checkpoints to check concepts, leveraging the generated structures and details at different stages.
@@
At each checkpoint, a Generation Check Block is utilized to check the presence of concepts. To effortlessly carry out the check of diverse concepts, we opt to incorporate a pre-trained Large Language-Vision Model (LLVM) [9] rather than training detection models for this task. It receives the predicted final images $\hat{x}_{0}$ as input and processes a predefined query, inquiring whether the images contain the target concepts. To format its outputs, we leverage in-context learning [8] in the formulation of the query. Please refer to Appendix A for the specific query we design.
@@ 4.3 Concept Removal Attention
We propose Concept Removal Attention, designed to remove the concepts checked in the Generation Check Mechanism. It replaces the cross-attention layers in text-to-image diffusion models, thereby achieving the replacement of concept-related content through the generative capabilities of the models. Its framework is shown in Fig.4.
@@
For each concept, we define a negative prompt $p_{n}$ to guide the models in generating alternative content. The attention function $CRAttn(z_{t},p,p_{n})$ is defined as:
@@
where $K_{n}=W_{k}\tau_{\theta}(p_{n})$ and $V_{n}=W_{v}\tau_{\theta}(p_{n})$. The primary difference between Eq.4 and Eq.6 is that Eq.6 has $\alpha_{t}$ which plays the role of a temperature coefficient. The motivation for this modification comes from our observation that the similarity between features of concept-related content and negative prompts is low sometimes. It causes the features of the prompt beginning token to occupy a high proportion of the attention weights, leading to the failure of this attention calculation. Recall the Softmax function $Softmax(\alpha_{t}s)_{i}=\frac{\exp^{\alpha_{t}s_{i}}}{\sum_{k}\exp^{\alpha_{t}s_{k}}}=\frac{1}{\sum_{k}\exp^{\alpha_{t}(s_{k}-s_{i})}}$. When $\alpha_{t}<1$, the weights corresponding to the large components in $s$ will become small. We leverage it to overcome the guidance failure of negative prompts. When $t$ goes from $T$ to 00, we set $\alpha_{t}=0.5+0.5\frac{t_{i}-t}{t_{i}}(t_{i}\in\{t_{1},t_{2}\})$ as an increasing function in the range of $[0.5,1]$.
@@
where $\mathbb{I}(\cdot)$ is the indicator function, $K_{c}=W_{k}\tau_{\theta}(c)$, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the matrix indexing operation, $idx_{EOS}$ is the index of the prompt ending token, and $\kappa$ is a threshold. Note that $c$ here stands for target concepts rather than any input prompt. It stems from our discovery that, even without explicit guidance from concept words, the features of concept words are closer to those of related content than those of other content. $idx_{EOS}$ is used because concepts such as nudity and Monet’s style consist of multiple tokens. The embedding features at $idx_{EOS}$ can encode the overall semantics of concepts. $\kappa$ is determined adaptively by the mean of the softmax scores. To obtain an accurate $\mathcal{M}_{1}$, we average the corresponding $\mathcal{M}_{1}$ across all attention heads and use this averaged $\mathcal{M}_{1}$ for the attention calculations of all heads. We also find that $\mathcal{M}_{1}$ is less accurate for the shallowest and deepest layers in the noise predictors (such as U-Net [38] in Stable Diffusion [37]). We speculate that the reason may be the smaller receptive field of high-resolution features in the shallowest layers and the more coarse features in the deepest layer, which limits their ability to recognize concepts. Therefore, in the first down-sampling layer, the attention is not applied. In the last up-sampling layer and the deepest layer, $\mathcal{M}_{1}$ is the average of $\mathcal{M}_{1}$ in the preceding layers. Please refer to Appendix B for the details. In Concept Removal Attention-2 for $t_{2}$, $\mathcal{M}=\mathcal{M}_{2}=\mathbf{1}$ is a matrix of all ones, intended for high-intensity, global replacement.
@@ 5.1 Experimental Setting
Baselines. 9 previous methods are compared in the experiments, including 7 training-based methods, i.e. CA [22], ESD [12], RECE [14], SalUn [10], MACE [29], Receler [18], and LatentGuard [26], and 2 training-free methods, i.e. SLD [42] and SAFREE [52]. The strength level of the safety guidance in SLD is set to Max. Please refer to Appendix C for the reproduction details.
@@
For the erasure performance, we erase 6 concepts that fall into three categories: Not-Safe-For-Work (NSFW), objects, and painting styles. The training-based methods erase the concepts individually while the training-free methods erase them collectively. Concept Ratio (%) measures the erasure performance under the user prompts and the adversarial prompts respectively. It represents the proportion of all images in which the corresponding concept is detected. GPT-4o [1] generates 100 user prompts for each concept except for the NSFW concepts. Please refer to Appendix D for the instruction for GPT-4o. For the NSFW concepts, the prompts are selected from Inappropriate Image Prompts (I2P) [42]. Compared with other prompts, the selected prompts enable SD v2.1 to generate images with the highest scores provided by the detectors. Based on these user prompts, Ring-A-Bell [46] searches the adversarial prompts. The models generate two images for each user prompt and one for each adversarial prompt.
@@
Implementations. The used LLVM is LLaVa-OneVision-Qwen2-7B [23], a recent model with state-of-the-art performance on various benchmarks. Other popular LLVMs can also achieve similar results in our experiments. The checkpoints $t_{1}$ and $t_{2}$ are set to 40 and 20 respectively. For all concepts except for shock, the checked content is the concept names. Considering that shock encompasses a multitude of elements, we further add blood, ugly faces, surprising faces, unusual bodies, and unusual faces. The compared methods have also embraced this supplement. The pre-defined negative prompts are listed in Appendix E.
@@ 5.3.1 Checkpoints
Table 4: Results with negative prompts (NP).
@@ 5.3.2 Concept Removal Attention
The impact of negative prompts. For erasing nudity, the used negative prompt is “Covered from neck to toe in clothing”. More prompts are evaluated. They include: ❶“dressed person”, ❷“person in clothes”, and ❸“Covered in clothing”. The results are shown in Tab.5.3.1. It confirms that our method is effective for different negative prompts.
@@ 5.3.4 Time Efficiency
Under the same implementation, our method is 40.7% faster than SAFREE in total runtime and 10.9% faster than SLD in generation time. The detailed results are presented in Appendix I. The efficiency bottleneck of our method is LLVM. Since it only acts as a binary concept classifier, training a tiny model can achieve a similar effect. Its smaller parameters as well as the application of acceleration methods such as quantization and pruning can improve checking efficiency. We leave the experiments for the future.
@@ 6 Conclusion
In this paper, we present Concept Corrector for erasing concepts in text-to-image diffusion models. By leveraging the generation mechanism of diffusion models, it checks intermediate-generated images and corrects concept-related content during the diffusion process. The experiments on various concepts demonstrate that our method can effectively erase these concepts, outperforming previous methods while maintaining the alignment between the images and concept-irrelevant prompts.
@@
    Query in Generation Check Mechanism.
@@
A Query in Generation Check Mechanism
We design the following query in the Generation Check Mechanism for the LLVM.
B Visualization of $\mathcal{M}_{1}$

### v2

@@ Abstract
Text-to-image diffusion models have demonstrated the underlying risk of generating various unwanted content, such as sexual elements. To address this issue, the task of concept erasure has been introduced, aiming to erase any undesired concepts that the models can generate. Previous methods, whether training-based or training-free, have primarily focused on the input side, i.e., texts. However, they often suffer from incomplete erasure due to limitations in the generalization from limited prompts to diverse image content. In this paper, motivated by the notion that concept erasure on the output side, i.e., generated images, may be more direct and effective, we propose Concept Corrector. It checks target concepts based on visual features provided by final generated images predicted at certain time steps. Further, it incorporates Concept Removal Attention to erase generated concept features. It overcomes the limitations of existing methods, which are either unable to remove the concept features that have been generated in images or rely on the assumption that the related concept words are contained in input prompts. In the whole pipeline, our method changes no model parameters and only requires a given target concept as well as the corresponding replacement content, which is easy to implement. To the best of our knowledge, this is the first erasure method based on intermediate-generated images, achieving the ability to erase concepts on the fly. The experiments on various concepts demonstrate its impressive erasure performance. Code.
Figure 1: We present Concept Corrector, a training-free method to erase concepts based on intermediate-generated images during the text-to-image diffusion process. It integrates the Generation Check Mechanism alongside Concept Removal Attention, to first check target concepts during generation and then erase them in the subsequent generations. Please refer to Appendix J for more visualizations.
@@ 1 Introduction
The task of concept erasure has been proposed to address this issue [42, 22, 11]. Its goal is to prevent models from generating content related to certain concepts. Intuitively speaking, if the generation path from the textual inputs to the target visual concepts can be cut off, the resultant images, being guided by those texts, would naturally exclude the presence of these concepts. Researchers have explored various methods based on this idea, which can be categorized into training-based and training-free methods. Within the area of training-based methods, CA [22] and ESD [11] fine-tune the generation distributions conditioned on the texts containing target concepts. Instead of gathering texts, Bui et al. [2] and Meng et al. [30] incorporate a learnable prompt in the training and other works [34, 21, 55, 18] introduce adversarial training to improve the robustness of concept erasure. UCE [12], MACE [29], and RECE [13] edit cross-attention weights by aligning the keys and values of target concepts to others. There are also some methods [9, 49, 4] ablating model parameters based on their sensitivity to related prompts. While these methods fine-tune diffusion models, Latent Guard [26] and GuardT2I [51] propose to fine-tune the text encoders. Within the area of training-free methods, Negative Guidance [37] and SLD [42] use concept texts to steer generation in the opposite direction. SAFREE [52] maps the prompt embeddings to those of target concepts and removes the corresponding components.
@@
While most studies explore text-based concept erasure methods, little attention has been paid to accomplishing this task based on the output of generative models, i.e. generated images. Text-based erasure methods often face the challenge of prompt generalization [46, 56, 50, 6]. It arises due to the difficulty in using prompts, which are collected or learned by these methods, to comprehensively cover the diverse image content associated with target concepts. Consequently, the effectiveness of the erasure is limited. Consider a simple example. The prompt “A woman looking seductive” does not explicitly convey the meaning of nudity. However, as demonstrated in Fig.1, it can prompt the erased models to generate an image of a naked woman. If the focus is shifted to images, the target of erasure becomes more directly addressed, potentially leading to a more effective erasure performance. A common practice is to incorporate a safety checker following the generation process to filter out unwanted content like in Stable Diffusion [37], which works well but cannot correct generated content to obtain images like text-based erasure methods.
In this paper, we aim to actively intervene intermediate images in the generation to reliably erase target concepts from final images. Using the diffusion generation theory, we note that final images predicted at certain time steps present enough structure and detail features for a detector to check concepts. Once concepts are detected, we consider how to erase them. Existing solutions like prompt editing [14] and negative guidance [15, 42] cannot remove the concept features from images, causing target concepts to still exist. Some methods like Receler [18] suppress concept features within the attention layers. However, they rely on concept words in input prompts, which are difficult to anticipate in advance or accurately capture during inference due to the diversity and implicitness of languages. Thus, we propose Concept Removal Attention, a variation of the cross attention mechanism. It erases generated concept features by giving the names of target concepts and negative concepts and perturbing generated features. These efforts form our method Concept Corrector.
@@
Compared with previous methods, our method has significant advantages in terms of reliability. Firstly, it checks concepts in images, thereby providing a more direct assessment of concepts. Secondly, it does not require input prompts to contain explicit concept words, solving the problem of prompt generalization. Moreover, all parameters remain unchanged, which protects model knowledge. In the experiments, we evaluate the erasure performance using user prompts and adversarial prompts. The evaluated concepts include Not-Safe-For-Work, objects, and painting styles. As shown in Fig.1, Tab.1 and Tab.2, our method achieves impressive erasure performance, significantly reducing the generation of most concepts to within 5% while other methods are still far from this level of performance. The contributions of this paper can be outlined as follows.
@@
  • We carefully analyze the feasibility of using intermediate images in the generation for checking target concepts.
  • We propose Concept Removal Attention to erase concept features in the generation. It changes no model parameters and only requires the names of target concepts and negative concepts, making erasure easy to implement.
@@
  • The above explorations forms Concept Corrector, a straightforward but effective method to erase concepts on the fly. To our knowledge, this is the first work to achieve erasure based on intermediate-generated images.
@@
  • The experiments and visualizations demonstrate the impressive effectiveness of our method. A series of ablation experiments are conducted to discuss each component.
@@ 4.2 Generation Check Mechanism
As mentioned in Sec.3.1, the final generation results can be predicted at intermediate time steps. We highlight that they provide rich visual features to check concepts during the generation. Some previous studies [7, 19] explore the generative traits of different diffusion stages. A common observation is that diffusion models initially generate global structures and then refine these with local details as the diffusion progresses. It inspires us to set two checkpoints to check concepts, leveraging the generated structures and details at different stages.
@@
At each checkpoint, we set a Generation Check Block respectively. It receives the predicted final images $\hat{x}_{0}$ as input and uses a detector to decide whether they contain any target concepts. Please see Sec.5.1 for the implementation.
@@ 4.3 Concept Removal Attention
Conditioned that there are target concepts in intermediate images, we further consider how to erase them. Editing prompts [14] or guiding generation negatively cannot remove features that already exist in images, leading to concepts still being present in final outputs. We provide the experiments and discussions in Appendix A.1. Some methods like Receler [18] suppress the features by concept words in input prompts, but the concept words we define may not appear in input prompts. To bridge this gap, we propose Concept Removal Attention, as shown in Fig.4.
@@
For each target concept, we define a negative concept $p_{n}$ to guide the models in generating alternative content. The attention function $CRAttn(z_{t},p,p_{n})$ is defined as:
@@
where $K_{n}=W_{k}\tau_{\theta}(p_{n})$ and $V_{n}=W_{v}\tau_{\theta}(p_{n})$. The primary difference between Eq.4 and Eq.6 is that Eq.6 has $\alpha_{t}$ which plays the role of a temperature coefficient. The motivation for this modification comes from our observation that the similarity between features of concept-related content and negative concepts is low sometimes. It causes the features of the prompt beginning token to occupy a high proportion of the attention weights, leading to the failure of this attention calculation. Recall the Softmax function $Softmax(\alpha_{t}s)_{i}=\frac{\exp^{\alpha_{t}s_{i}}}{\sum_{k}\exp^{\alpha_{t}s_{k}}}=\frac{1}{\sum_{k}\exp^{\alpha_{t}(s_{k}-s_{i})}}$. When $\alpha_{t}<1$, the weights corresponding to the large components in $s$ will become small. We leverage it to overcome the guidance failure of negative prompts. When $t$ goes from $T$ to 00, we set $\alpha_{t}=0.5+0.5\frac{t_{i}-t}{t_{i}}(t_{i}\in\{t_{1},t_{2}\})$ as an increasing function in the range of $[0.5,1]$.
@@
where $\mathbb{I}(\cdot)$ is the indicator function, $K_{c}=W_{k}\tau_{\theta}(c)$, [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] denotes the matrix indexing operation, $idx_{EOS}$ is the index of the prompt ending token, and $\kappa$ is a threshold. Note that $c$ here stands for target concepts rather than any input prompt. It stems from our discovery that, even without explicit guidance from concept words, the features of concept words are closer to those of related content than those of other content. $idx_{EOS}$ is used because concepts such as nudity and Monet’s style consist of multiple tokens. The embedding features at $idx_{EOS}$ can encode the overall semantics of concepts. $\kappa$ is determined adaptively by the mean of the softmax scores. To obtain an accurate $\mathcal{M}_{1}$, we average the corresponding $\mathcal{M}_{1}$ across all attention heads and use this averaged $\mathcal{M}_{1}$ for the attention calculations of all heads. We also find that $\mathcal{M}_{1}$ is less accurate for the shallowest and deepest layers in the noise predictors (such as U-Net [38] in Stable Diffusion [37]). We speculate that the reason may be the smaller receptive field of high-resolution features in the shallowest layers and the more coarse features in the deepest layer, which limits their ability to recognize concepts. Therefore, in the first down-sampling layer, the attention is not applied. In the last up-sampling layer and the deepest layer, $\mathcal{M}_{1}$ is the average of $\mathcal{M}_{1}$ in the preceding layers. Please refer to Appendix A.2 for the details. In Concept Removal Attention-2 for $t_{2}$, $\mathcal{M}=\mathcal{M}_{2}=\mathbf{1}$ is a matrix of all ones, intended for high-intensity, global replacement.
@@ 5.1 Experimental Setting
Baselines. 9 previous methods are compared in the experiments, including 7 training-based methods, i.e. CA [22], ESD [11], RECE [13], SalUn [9], MACE [29], Receler [18], and LatentGuard [26], and 2 training-free methods, i.e. SLD [42] and SAFREE [52]. The strength level of the safety guidance in SLD is set to Max. Please refer to Appendix B for the reproduction details.
@@
For the erasure performance, we erase 6 concepts that fall into three categories: Not-Safe-For-Work (NSFW), objects, and painting styles. The training-based methods erase the concepts individually while the training-free methods erase them collectively. Concept Ratio (%) measures the erasure performance under the user prompts and the adversarial prompts respectively. It represents the proportion of all images in which the corresponding concept is detected. GPT-4o [1] generates 100 user prompts for each concept except for the NSFW concepts. Please refer to Appendix C for the instruction for GPT-4o. For the NSFW concepts, the prompts are selected from Inappropriate Image Prompts (I2P) [42]. Compared with other prompts, the selected prompts enable SD v2.1 to generate images with the highest scores provided by the detectors. Based on these user prompts, Ring-A-Bell [46] searches the adversarial prompts. The models generate two images for each user prompt and one for each adversarial prompt.
@@
Implementations. To effortlessly check diverse concepts, we opt to incorporate a pre-trained Vision-Language Model (VLM) [8] rather than training extra detection models. Specifically, the used VLM is LLaVa-OneVision-Qwen2-7B [23], a recent model with state-of-the-art performance on various benchmarks. Other popular VLMs can also achieve similar results in our experiments. The designed query for the VLM can be found in Appendix D. The checkpoints $t_{1}$ and $t_{2}$ are set to 40 and 20 respectively. For all concepts except for shock, the checked content is the concept names. Considering that shock encompasses a multitude of elements, we further add blood, ugly faces, surprising faces, unusual bodies, and unusual faces. The compared methods have also embraced this supplement. The pre-defined negative concepts are listed in Appendix E.
@@ 5.3.1 Checkpoints
Table 4: Results with negative concepts (NC).
@@ 5.3.2 Concept Removal Attention
The impact of negative concept descriptions. For erasing nudity, we use “Covered from neck to toe in clothing” to describe the negative concept. More descriptions are evaluated. They include: ❶“dressed person”, ❷“person in clothes”, and ❸“Covered in clothing”. The results are shown in Tab.5.3.1. It confirms that our method is robust to various descriptions of a negative concept.
@@ 5.3.4 Time Efficiency
Under the same implementation, our method is 10.9% faster than SLD and 75.4% faster than SAFREE in the generation time. It should be noticed that simple concept-specific detectors rather than the VLM can also achieve a similar binary check performance but improve the efficiency of the pipeline significantly, as demonstrated by our extended experiments. Please refer to Appendix I for these details.
@@ 6 Conclusion
In this paper, we introduce and validate the feasibility of erasing concepts based on intermediate generated images rather than input prompts, a straightforward yet under-explored approach that is neglected in existing studies. Capitalizing on this insight, we propose Concept Corrector for erasing concepts on the fly without changing any parameters. Quantitative experiments coupled with visualizations demonstrate that it can reliably erase unwanted concepts while aligning images and non-targeted textual descriptors.
@@
    Concept Removal Attention.
@@
  11. K.
    Limitations.
A Concept Removal Attention
A.1 Results of Other Alternatives
In the main paper, we highlight that existing methods, such as prompt editing and negative guidance, cannot remove the generated concept-related features from intermediate images, leading to concepts still being present in the final outputs. In this subsection, we provide the specific experimental results. Specifically, without changing our pipeline, we replace our proposed Concept Removal Attention with Prompt-to-Prompt [14] and Negative Guidance. The target concept is nudity. For Prompt-to-Prompt, we use GPT-4o to edit the user prompts into the ones without the sexual meaning by instructing it to add, replace, or remove some words. For Negative Guidance, we set “nudity, nude, naked, sexual, exposed, unclothed” as the negative prompts. Moreover, we also compare the performance of the two methods without being integrated into our pipeline. The evaluation protocol follows the one reported in the main paper.
[…]
SD v2.1 61.5
Negative Guidance 18.5
Prompt-to-prompt 37.5
Ours (+ Negative Guidance) 48.5
Ours (+ Prompt-to-Prompt) 50.0
Ours 4.0
Table S9: The erasure results of Negative Guidance, Prompt-to-Prompt, and the methods which are integrated into our proposed pipeline. The concept is nudity, and the Concept Ratio is reported.
Figure S6: The visualizations of the images with the prompts containing the concept of nudity. NG: Negative Guidance. P2P: Prompt-to-Prompt. Ours (+ NG): Replace Concept Removal Attention with NG in our proposed method. Ours (+ P2P): Replace Concept Removal Attention with P2P in our proposed method.
Tab.S9 presents the results. From the table, we can see that when we integrate Negative Guidance and Prompt-to-Prompt into our proposed generation pipeline, the nudity ratio of the generated images only drops slightly compared with the original model. Fig.S6 gives some examples. The generated images after their erasure, i.e. Ours (+NG) and Ours (+P2P), are highly similar to the original images. In addition, the prompts only contain the implicit words corresponding to nudity, such as bathhub, underwear, and bather, leading to the difficulty to erase nudity for Prompt-to-Prompt. Even if we apply Negative Guidance and Prompt-to-Prompt from the very beginning of generation, the concept of nudity is still not removed. On the contrary, our proposed method erase the concept successfully, achieving the least nudity ratio among these alternatives.
A.2 Visualization of $\mathcal{M}_{1}$